A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a long-term government and corporate bond fund, and the third is a T-bill money market fund that yields a sure rate of 4.4%. The probability distributions of the risky funds are:

Expected Return Standard Deviation
Stock fund (S) 14% 34%
Bond fund (B) 5% 28%

The correlation between the fund returns is 0.0214.

What is the expected return and standard deviation for the minimum-variance portfolio of the two risky funds? (Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

Weight of (S) = ((Standard Deviation of B)^2 -correlation * Standard Deviation of S * Standard Deviation of B)/((Standard Deviation of S)^2 + (Standard Deviation of B)^2 - 2 * correlation * Standard Deviation of S * Standard Deviation of B)
=((28%^2)-0.0214*34%*28%)/((34%^2)+(28%^2)-2*0.0214*34%*28%)=40.2066832121068%
Weight of B =1-40.2066832121068%=59.7933167878932%

Expected Return of min variance portfolio =40.2066832121068%*14%+59.7933167878932%*5%=9.62%
Standard Deviation = ((Weight of S * Standard Deviation of S)^2 + (weight of B * standard Deviation of B)^2 + 2* Weight of S * Standard Deviation of S * weight of B * standard Deviation of B * correlation)^0.5=((40.2066832121068%*34%)^2+(59.7933167878932%*28%)^2+2*40.2066832121068%*59.7933167878932%*34%*28%*0.0214)^0.5=21.84%